# Learning Task-adaptive Quasi-stiffness Control for a Powered Transfemoral Prosthesis

### v1

@@ Abstract
While significant advancements have been made in the mechanical and task-specific controller designs of powered transfemoral prostheses, developing a task-adaptive control framework that generalizes across various locomotion modes and terrain conditions remains an open problem. This study proposes a task-adaptive learning quasi-stiffness control framework for powered prostheses that generalizes across tasks, including the torque-angle relationship reconstruction part and the quasi-stiffness controller design part. Quasi-stiffness is defined as the slope of the human joint’s torque-angle relationship. To accurately obtain the torque-angle relationship in a new task, a Gaussian Process Regression (GPR) model is introduced to predict the target features of the human joint’s angle and torque in the task. Then a Kernelized Movement Primitives (KMP) is employed to reconstruct the torque-angle relationship of a new task from multiple human demonstrations and estimated target features. Based on the torque-angle relationship of the new task, a quasi-stiffness control approach is designed for a powered prosthesis. Finally, the proposed framework is validated through practical examples, including varying speed and incline walking tasks. The proposed framework has the potential to expand to variable walking tasks in daily life for the transfemoral amputees.
I Introduction
@@
Transfemoral amputees suffer ambulation difficulties when performing demanding locomotion modes, such as climbing ramps and stairs. Powered transfemoral prostheses have the potential to restore the lost function of the joints and enhance the mobility of the amputees [1, 2]. While significant progress has been made in both the mechanical designs [3, 4, 5, 6, 7, 8, 9, 10] and controller designs [11, 12, 13, 14, 15, 16, 17] of powered prostheses over the past decades, developing a prosthesis control framework that generalizes across a variety of locomotion modes and terrain conditions remains a challenge [2].
@@
Impedance-based control is the most commonly used approach in powered prostheses because of its ability to reproduce natural walking. These methods segment the gait cycle into several sub-phases, each performing an impedance control law. Previous researchers manually tuned the impedance parameters for each task, such as flat walking [3], ramp climbing [18], and stair climbing [19]. However, the heuristic parameter tuning procedure for a specific task takes at least several hours [20]. As shown in Fig. 1, transfemoral amputees may perform different tasks (locomotion modes and terrain conditions) in daily walking. The time required for tuning a powered prosthesis for each task and each patient presents a major barrier to its clinical feasibility [21]. To reduce the tuning time, reinforcement learning (RL) is implemented in [14, 22] to automatically tune the impedance parameters. Nevertheless, these approaches are limited to flat walking task and knee joint only at present. Recent studies have achieved prostheses handling variable-task walking. Data-driven, impedance-based controllers are respectively designed for varying speeds and inclines walking [20], stair climbing [23], and sit/stand [24]. An adaptive stair controller enables transfemoral amputee patients to climb stairs of varying stair heights [25]. These controllers are still required to extend to a unified framework for variable-task walking. To date, developing powered transfemoral prostheses controllers that generalize across tasks remains an open problem.
@@
Quasi-stiffness control is a subset of the impedance-based control strategies[2]. This strategy determines the impedance parameters of each sub-phase from the torque-angle relationship of human’s lower limb joints. The slope of the torque-angle relationship is called ”quasi-stiffness” [2, 26]. The desired prosthesis torque is then determined by the quasi-stiffness and equilibrium angle as . The torque–angle relationships can be modified for quasi-stiffness control of powered prostheses in different tasks. For example, the stair controller modulates the torque-angle relationship based on the stair height for stair climbing tasks. In addition, it is reported that quasi-stiffness control strategies have the potential to improve gait symmetry and achieve speed-adaption without subject-specific or speed-specific tuning [21]. Inspired by the quasi-stiffness control strategy, this study proposes a learning task-adaptive quasi-stiffness control framework that generalizes across a variety of tasks. Two issues are presented and need to be resolved. The first one is to automatically construct joints’ torque-angle relationship in a new task. The second one is to design quasi-stiffness controllers based on the generated torque-angle relationship.
@@
The torque-angle relationship is determined by the joint angle and torque and is widely used in the design and control of powered exoskeleton [27, 28] and prosthesis [21, 25]. The peak values (named as target features in this study) in the profiles of joint kinematics and kinetics vary with tasks and are often used to represent the joint kinematics and kinetics during walking tasks [14, 22]. However, most torque-angle relationships and target features are obtained in the laboratory by a motion capture system. It is impossible to collect joint kinematics and kinetics for all tasks. This study proposed an imitation learning-based method that can reconstruct the torque-angle relationships of new tasks by learning from multiple demonstrated torque-angle relationships. For example, reconstructing a relationship at a new walking speed from learning the existing demonstrated relationship at different speeds (or a new relationship at a new ramp inclination, etc.). To reconstruct the torque-angle relationship of a new task, the Gaussian Process Regression (GPR) [29, 30] is introduced to predict the target features of the joint angles and torques based on the demonstrated features and task variables (such as speed and terrain parameters). Furthermore, the new task torque-angle relationship can be reconstructed by the target features and human-demonstrated torque-angle relationships. While many fitting methods can use these target features for torque-angle relationship reconstruction, such as polynomial interpolation [31] and Fourier Series fitting [32], these methods consider only trajectory fitting instead of the multiple demonstrated relationships. Instead, learning from multiple demonstrations methods such as probabilistic trajectory methods can be introduced. Gaussian Mixture Model/Regression (GMM-GMR) [33, 34] was applied to reproduce joint angles and learning trajectory distribution through learning from the demonstrated database. However, this method makes it difficult to modify the learned trajectories to fulfill additional new constraints, such as passing through specific target features. A Kernelized Movement Primitives (KMP) [35] method was presented that allows the reproduced learning from demonstration trajectories passing through the specific via-points. Inspired by this, KMP is employed to construct the new torque-angle relationship from multiple demonstrated relationships and target features. After obtaining the torque-angle relationship, the quasi-stuffiness control strategy is introduced to design a task-adaptive controller that generalizes across tasks.
@@
A learning-based torque-angle relationship construction method is proposed for various tasks.
@@
A quasi-stiffness control strategy based on the learned torque-angle relationship is presented for powered transfemoral prostheses.
3.
Examples of varying speed and varying incline walking using the proposed quasi-stiffness control framework are conducted and validated through walking experiments with a transfemoral amputee.
@@ II Method
As shown in Fig. 2, the proposed task-adaptive learning quasi-stiffness control framework consists of a learning from human demonstrations part and a quasi-stiffness control part. The learning-based part is composed of two models: a GPR-based target features prediction model and a KMP-based task-specific torque-angle relationship reconstruction model.
II-A Task Estimation
Our previous studies and some other studies presented various task estimation methods, such as vision-based methods [36, 37] for terrain condition estimation and IMU (inertial measurement unit)-based methods [20, 38, 39] for speed estimation. These methods can be applied to estimate the task parameter (where denotes the walking speed, represents the terrain parameters) for the proposed framework.
II-B GPR-based Target Features Prediction
An open-source dataset [40] is implemented as the human demonstrations, which contains tasks including multiple locomotion modes (flat walking, ramp climbing, and stair climbing) and terrain conditions (walking speeds, ramp inclinations, and stair heights). To obtain the new task-specific torque-angle relationship, this study introduces GPR-based learning from the demonstration method for target features prediction, as shown in Fig. 2(b). Considering the limited human demonstration size, GPR, a nonparametric kernel-based probabilistic model, adeptly balances fitting accuracy with model complexity, mitigating the risk of overfitting in small-size training samples. Assume there are demonstrations and define the train data as
where denotes the number of the features; is the task parameter of the -th demonstration (-th task). is the output features (joint angles/torques) of the -th demonstration. For the task , the GPR [29, 30] is intended to transform the input task vector into a output value by , where represents the Gaussian noise satisfying . The training dataset can be described by the GPR model as
@@ II-B GPR-based Target Features Prediction
II-C KMP-based Torque-angle relationship Reconstruction
After obtaining the kinematics and kinetics target features of the new task, the KMP is employed to reconstruct the torque-angle relationship with the target features of the new task, as shown in Fig. 2(c). Denote the set of demonstrated torque-anlge relationship by where represents the joint kinematics and is the joint kinetics. and denote the length of the torque-angle relationship and the number of the demonstrations. GMM [34] is introduced to encode the demonstrated dataset and estimate the relationship probability distribution
@@ II-C KMP-based Torque-angle relationship Reconstruction
where and represent the reference torque and angle respectively, and denote mean and variance respectively. As mentioned before, we expect that the new torque-angle relationship can pass through the target features of the new task. Therefore, KMP is employed to learn from demonstrations and fulfill the additional constraints. Considering a parametric torque-angle relationship can be formulated by KMP as
@@ II-D Quasi-stiffness Estimation and Controller Design
where denotes the quasi-stiffness, represents the equilibrium angle. It is noted that the torque-angle relationship during the stance phase is approximately linear [26, 43]. Inspired by this, as shown in Fig. 2(d), this study applies a linear regression to estimate the quasi-stiffness of the torque-angle relationship, and the equilibrium angle is equal to the value of the intersection point between the regression equation and the x-axis. For the prosthesis control, we divide the gait phase into four sub-phases: stance flexion, stance extension, swing flexion, and swing extension. The sub-phase switching follows a finite-state machine (FSM), as shown in Fig. 3. For the stance phase, the torque-angle relationship is approximately linear, which can be approximated by a linear equation. While the swing phase is much more complex than the stance phase [43], we choose two linear parts that can achieve swing flexion and extension and use linear functions to approximate these two parts.
III Results
III-A Experimental Setup
@@
The powered transfemoral prosthesis used in this study is shown in Fig. 6(a). Both knee and ankle joints are driven by frameless motors (ILM motor kit, RoboDrive, Seefeld, Germany) with a 50:1 harmonic reducer (CSD-25-50-2A-GR, Nagano Prefecture, Japan). The frameless motors are actuated by drives (G-MOLTWIR50/100SEO, Elmo Motion Control, Petah Tikva, Israel). A 6-axis force/torque sensor (Sunrise Instruments, Nanning, China) is employed to measure the (ground reaction forces) GRFs.
A transfemoral amputee patient is invited to participate in the experiments to evaluate the proposed learning quasi-stiffness control framework. The participant provided informed consent, and the Southern University of Science and Technology Institutional Review Board approved the experiments.
III-B Experimental Results of the Proposed Framework
@@
Due to the space limitations, this study presents an example of varying walking speed tasks to show how the learning quasi-stiffness control framework works. The learning process of the GPR-based target feature estimation and KMP-based torque-angle relationship construction is offline on the joint kinematics and kinetics obtained from the open-source dataset [40].
@@
1) Target Features Estimation: Fig. 2(b) shows the experimental results of target features prediction, where the blue dots denote the demonstrated target features and the purple dots represent the predicted target features of a new task (new speed). The GPR-based target features prediction model is imported into the embedded platform of the powered prosthesis and used for feature prediction of new tasks.
@@
2) Torque-angle relationship Reconstruction: Fig. 4 shows the torque-angle relationship learning and reconstruction in a new task. The distribution of the demonstrated torque-angle relationship can be learned by GMM. Based on the learned GMM, the reference torque-angle relationship can be reproduced by GMR. As shown in Fig. 4(a), the blue torque-angle curves denote the human demonstrations of different speed walking tasks, and the yellow torque-angle curve represents the reference torque-angle relationship reproduced by GMR. As shown in Fig. 4(b), the yellow dashed curve represents a torque-angle relationship reconstructed by the demonstrated torque-angle relationships (blue curves) and target features (purple dots) in a new task.
@@
3) Quasi-stiffness Estimation: Fig. 5 shows using linear functions to approximate several parts of the knee and ankle torque-angle relationships in the varying speed walking tasks. Table I shows the estimated stiffness and equilibrium angles from the torque-angle relationships of different speed tasks and these impedance parameters are used in the powered transfemoral prosthesis experiments.
@@
4) Performance of the Proposed Method in Experiments: Fig. 6(b) shows a series of photos of a transfemoral amputee performing varying speed and inclines walking tasks. Fig. 7 presents that the kinematic and kinetics trajectories produced by the proposed method are similar to the able-bodied references. The powered prosthetic joint trajectories show strong similarities to the intact joint trajectories, which also indicate that the quai-stiffness can improve gait symmetry. Fig. 8 shows the root mean squared error (RMSE) in the powered prosthetic joint trajectories relative to able-bodied trajectories and intact leg trajectories during the walking experiments. The lower RMSE of the observed kinematics with respect to the intact leg kinematics indicates that the proposed quasi-stiffness control achieves better joint kinematic symmetry.
| Task (m/s) | Joint | Quasi-stiffness (Nm/deg) | Equilibrium Angle (deg) | ||||||
| 0.4 | knee | 4.13 | 0.59 | 0.08 | 0.19 | 12 | 22 | 30 | 15 |
| ankle | 0.82 | 6.14 | 0.53 | 0.59 | -5 | -2 | 8 | -8 | |
| 0.6 | knee | 5.68 | 0.70 | 0.17 | 0.26 | 12 | 21 | 57 | 17 |
| ankle | 0.76 | 5.05 | 0.53 | 0.64 | -5 | 0 | 8 | -8 | |
| 0.8 | knee | 5.79 | 0.79 | 0.27 | 0.45 | 13 | 22 | 69 | 18 |
| ankle | 0.78 | 4.23 | 0.45 | 0.69 | -6 | -4 | 9 | -7 | |
| 1.0 | knee | 5.85 | 0.85 | 0.36 | 0.52 | 13 | 22 | 76 | 20 |
| ankle | 1.03 | 3.89 | 0.36 | 0.72 | -5 | -7 | 10 | -8 | |

### v2

# Tuning-free Quasi-stiffness Control Framework of a Powered Transfemoral Prosthesis for Task-adaptive Walking

@@ Abstract
Impedance-based control represents a prevalent strategy in the development of powered transfemoral prostheses. However, creating a task-adaptive, tuning-free controller that effectively generalizes across diverse locomotion modes and terrain conditions continues to be a significant challenge. This letter proposes a tuning-free and task-adaptive quasi-stiffness control framework for powered prostheses that generalizes across various walking tasks, including the torque-angle relationship reconstruction part and the quasi-stiffness controller design part. A Gaussian Process Regression (GPR) model is introduced to predict the target features of the human joint’s angle and torque in a new task. Subsequently, a Kernelized Movement Primitives (KMP) is employed to reconstruct the torque-angle relationship of the new task from multiple human reference trajectories and estimated target features. Based on the torque-angle relationship of the new task, a quasi-stiffness control approach is designed for a powered prosthesis. Finally, the proposed framework is validated through practical examples, including varying speeds and inclines walking tasks. Notably, the proposed framework not only aligns with but frequently surpasses the performance of a benchmark finite state machine impedance controller (FSMIC) without necessitating manual impedance tuning and has the potential to expand to variable walking tasks in daily life for the transfemoral amputees.
Index Terms:
Tuning-free, task-adaptive, prosthesisI Introduction
@@
Individuals with transfemoral amputations suffer difficulties in ambulation, particularly during demanding locomotion activities such as ascending ramps and stairs. Powered transfemoral prostheses hold promise in restoring the lost joint function and significantly enhancing the mobility of these individuals [1, 2]. While significant progress has been made in both the mechanical designs [3, 4, 5, 6, 7, 8] and controller designs [9, 10, 11, 12, 13, 14, 15, 16] of powered prostheses over the past y, developing a prosthesis control framework that generalizes across various locomotion modes and terrain conditions remains a challenge [2].
@@
Impedance-based control is the most commonly used approach in powered prostheses because of its ability to reproduce natural walking. These methods segment the gait cycle into several sub-phases, each performing an impedance control law. Previous researchers manually tuned the impedance parameters for each task, such as flat walking [3], ramp climbing [17], and stair climbing [18]. However, the heuristic parameter tuning procedure for a specific task takes at least several hours [19]. Transfemoral amputees may perform different tasks (locomotion modes and terrain conditions) in daily walking. The time required to tune a powered prosthesis for each task and each patient presents a significant barrier to its clinical feasibility [20]. To reduce the tuning time, reinforcement learning (RL) is implemented in [12, 21] to tune the impedance parameters automatically. Nevertheless, these approaches are limited to flat walking task and knee joint only at present. Recent studies have achieved prostheses handling variable-task walking. Data-driven, impedance-based controllers are respectively designed for varying speeds and inclines walking [19], stair climbing [22], and sitting/standing [23]. An adaptive stair controller enables transfemoral amputee patients to climb stairs of varying heights [24]. These controllers are still required to extend to a unified framework for variable-task walking. To date, developing powered transfemoral prostheses controllers that generalize across tasks remains an open problem.
@@
Quasi-stiffness control is a subset of the impedance-based control strategies [2]. While traditional impedance control requires tuning numerous impedance parameters to generate biologically joint torques, this quasi-stiffness strategy determines the impedance parameters of each sub-phase from the torque-angle relationship of human lower limb joints. The slope of the torque-angle relationship is called ”quasi-stiffness” [2, 25]. The desired prosthesis torque is then determined by the quasi-stiffness and equilibrium angle as . The torque–angle relationships can be modified for quasi-stiffness control of powered prostheses in different tasks. For example, the stair controller modulates the torque-angle relationship based on the stair height for stair climbing tasks [24]. While quasi-stiffness control strategies have been proven can improve gait symmetry and achieve speed-adaption without speed-specific tuning [20], how to adapt to other tasks automatically is still a challenge.
@@
Inspired by the previous research on quasi-stiffness control strategies, this study proposes a tuning-free quasi-stiffness control framework that generalizes across a variety of tasks. Firstly, a novel task-adaptive torque-angle relationship reconstruction method is proposed for autonomously generating the torque-angle relationship of a new task. The torque-angle relationship is constructed by the the joint angle and torque and its profile depends strongly on the peak values of the joint angle and torque. The peak values (named as target features in this study) in the profiles of joint angles and torques vary with tasks and are often used to represent the joint kinematics and kinetics during walking tasks [12, 21]. However, most torque-angle relationships and target features are obtained in the laboratory by a motion capture system. It is only possible to collect joint kinematics and kinetics for some tasks. This study proposes an imitation learning-based method that can reconstruct the torque-angle relationships of new tasks by learning from multiple reference torque-angle relationships. For example, reconstructing a relationship at a new walking speed from learning the existing reference relationships at different speeds (or a new relationship at a new ramp inclination, etc.). To reconstruct the torque-angle relationship of a new task, the Gaussian Process Regression (GPR) [26, 27] is introduced to predict the target features of the joint angles and torques based on the reference features and task variables (such as speeds and terrain parameters). Considering the limited human reference size, GPR, a nonparametric kernel-based probabilistic model, adeptly balances fitting accuracy with model complexity, mitigating the risk of over-fitting in small-size training samples. Furthermore, the torque-angle relationship of the new task can be reconstructed by the target features and human reference torque-angle relationships. While many fitting methods can use these target features for torque-angle relationship reconstruction, such as polynomial interpolation [28] and Fourier Series fitting [29], these methods consider only trajectory fitting instead of the multiple reference relationships. Instead, learning from multiple reference trajectory methods, such as probabilistic trajectory methods can be introduced. Gaussian Mixture Model/Regression (GMM/GMR) [30, 31] was applied to reproduce joint angles and learning trajectory distribution through learning from the reference database. However, this method makes it difficult to modify the learned trajectories to fulfill additional constraints, such as passing through specific target features. A Kernelized Movement Primitives (KMP) [32] method was presented that allows the reproduced learning from reference trajectories passing through the specific via-points. Inspired by this, KMP is employed to construct the new torque-angle relationship from multiple reference relationships and target features. Secondly, the quasi-stuffiness control strategy is introduced after obtaining the torque-angle relationship to design a task-adaptive controller that generalizes across tasks.
@@
A learning-based task-adaptive torque-angle relationship construction method is proposed for powered transfemoral prostheses, which can autonomously generate each task-specific torque-angle relationship during continuously varying tasks. A tuning-free quasi-stiffness control method based on the learned torque-angle relationship is presented for powered transfemoral prostheses.
@@
Experimental results of varying speed and varying incline walking tasks show that the proposed control framework reproduces biomimetic kinematics and kinetics similar to able-bodied references, meets the performance of the most widely used (one of the state of the art) FSMIC without requiring manual impedance tuning, and has the potential to enable autonomous task adaptation during continuously varying walking tasks.
@@ II Method
II-A Overview of the Framework
As shown in Fig. 1, the proposed tuning-free, task-adaptive quasi-stiffness control framework consists of a task estimation part, a torque-angle relationship reconstruction part, and a quasi-stiffness control part. The task estimation part perceives amputee locomotion information and terrain conditions. It provides task parameters ( where denotes walking speed and represents terrain parameters, such as inclination and stair height .) for the next part. Previous research, including our own, has explored a variety of task estimation methodologies, encompassing vision-based approaches [33, 34] for assessing terrain conditions and IMU (Inertial Measurement Unit)-based strategies [19, 35, 36] for estimating walking speed. These methodologies facilitate the estimation of task parameters .
The torque-angle relationship reconstruction part includes a GPR-based target features prediction model and a KMP-based task-specific torque-angle relationship reconstruction model. This part automatically constructs the joints’ torque-angle relationship in a new task. The quasi-stiffness control part determines the parameters and applies to a quasi-stiffness controller based on the generated torque-angle relationship. The details of these two parts are as follows.
II-B Target Features Prediction
An open-source dataset [37] is implemented as the human reference trajectories, which contains tasks including multiple locomotion modes (flat walking, ramp climbing, and stair climbing) and terrain conditions (walking speeds, ramp inclinations, and stair heights). To obtain the new task-specific torque-angle relationship, this study introduces GPR-based learning from the reference trajectory method for target features prediction, as shown in Fig. 1(b). Assume there are reference trajectories and define the train data as
where denotes the number of the features; is the task parameter of the -th reference trajectory (-th task). is the output features (joint angles/torques) of the -th reference trajectory. For the task , the GPR [26, 27] is intended to transform the input task vector into a output value by , where represents the Gaussian noise satisfying . The training dataset can be described by the GPR model as
@@ II-B Target Features Prediction
II-C Torque-angle Relationship Reconstruction
After obtaining the kinematics and kinetics target features of the new task, the KMP is employed to reconstruct the torque-angle relationship with the target features of the new task, as shown in Fig. 1(c). Denote the set of reference torque-angle relationship by where represents the joint kinematics and is the joint kinetics. and denote the length of the torque-angle relationship and the number of the reference trajectories. GMM [31] is introduced to encode the reference dataset and estimate the relationship probability distribution
@@ II-C Torque-angle Relationship Reconstruction
where and represent the reference torque and angle respectively, and denote mean and variance respectively. As mentioned before, we expect that the new torque-angle relationship can pass through the target features of the new task. Therefore, KMP is employed to learn from reference trajectores and fulfill the additional constraints. Considering a parametric torque-angle relationship can be formulated by KMP as
@@ II-D Quasi-stiffness Estimation and Controller Design
where denotes the quasi-stiffness, represents the equilibrium angle. For the prosthesis control, following the quasi-stiffness trends observed in able-bodied subjects [20], we divide the gait phase into four sub-phases: stance flexion, stance extension, swing flexion, and swing extension. The sub-phase switching follows a finite-state machine (FSM), as shown in Fig. 2. It is noted that the torque-angle relationship during the stance phase is approximately linear [25, 40]. Inspired by this, as shown in Fig. 1(d), this study applies a linear regression to estimate the quasi-stiffness of the torque-angle relationship, and the equilibrium angle is equal to the value of the intersection point between the regression equation and the x-axis. While the swing phase is much more complex than the stance phase [40], we choose two linear parts that can achieve swing flexion and extension and use linear functions to approximate these two parts.
III Experiments and Results
III-A Experimental Setup and Protocol
@@
1) Prosthesis Hardware: The powered transfemoral prosthesis used in this study is shown in Fig. 3. Both knee and ankle joints are driven by frameless motors (ILM motor kit, RoboDrive, Seefeld, Germany) with a 50:1 harmonic reducer (CSD-25-50-2A-GR, Nagano Prefecture, Japan). The frameless motors are actuated by drives (G-MOLTWIR50/100SEO, Elmo Motion Control, Petah Tikva, Israel). A 6-axis force/torque sensor (Sunrise Instruments, Nanning, China) is employed to measure the GRFs (ground reaction forces).
2) Experimental Protocol: A transfemoral amputee patient (44 years old, 60 kg, 170 cm, 24 years post-amputation) is invited to participate in the experiments to evaluate the proposed learning quasi-stiffness control framework. To investigate the task-adaptive performance of the proposed control framework, steady-state walking task trails including varying speeds (from 0.4 m/s to 1.0 m/s, m/s deviations) and varying inclines (from 2 deg to 8 deg, deg deviations) are conducted on an instrumented split-belt treadmill (Bertec, Columbus, OH, USA). The initial trials began with a period of acclimation in which the amputee participant first walked on the treadmill at a self-preferred speed (about 0.6 m/s) until feeling comfortable. After tuning and acclimation for a period of time, the participant walked on the treadmill again and performed each steady-state task trail for 30 seconds. The participant provided informed consent, and the Southern University of Science and Technology Institutional Review Board approved the experiments.
3) Benchmark FSMIC Setup: For comparison, this article introduces a benchmark FSMIC based on the existing studies [3, 17]. The selection of this FSMIC as a benchmark stems from its extensive usage, simplicity, and capacity to reproduce human-like kinematics and kinetics after appropriate impedance tuning. During the benchmark FSMIC tuning period, combined with feedback from both the participants and the prosthetist, the tuning process persisted until the authors, prosthetist, and participant expressed satisfaction with the attained natural gait.
III-B Experimental Results
@@
Due to the space limitations, this study presents an example of varying walking speed tasks and varying inclines walking tasks to show how the learning quasi-stiffness control framework works. The learning process of the GPR-based target feature estimation and KMP-based torque-angle relationship construction is offline on the joint kinematics and kinetics obtained from the open-source dataset [37].
@@
1) Target Features Estimation: Fig. 1(b) shows the experimental results of target features prediction, where the blue dots denote the reference target features and the purple dots represent the predicted target features of a new task (new speed). The GPR-based target features prediction model is imported into the embedded platform of the powered prosthesis and used for feature prediction of new tasks.
@@
2) Torque-angle Relationship Reconstruction: Fig. 4 shows the torque-angle relationship learning and reconstruction in a new task. The distribution of the human reference torque-angle relationship can be learned by GMM. Based on the learned GMM, the reference torque-angle relationship can be reproduced by GMR. As shown in Fig. 4(a), the blue torque-angle curves denote the human reference trajectoires of different speed walking tasks, and the yellow torque-angle curve represents the reference torque-angle relationship reproduced by GMR. As shown in Fig. 4(b), the yellow dashed curve represents a torque-angle relationship reconstructed by the human reference torque-angle relationships (blue curves) and target features (purple dots) in a new task.
@@
3) Quasi-stiffness Estimation: Fig. 5 shows using linear functions to approximate several parts of the knee and ankle torque-angle relationships in the varying speed walking tasks.
@@
4) Amputee Participant Experimental Results: Fig. 6 shows a series of gait trails of a transfemoral amputee performing varying speed and inclines walking tasks. Fig. 7 presents that the kinematic and kinetics trajectories produced by the proposed method are similar to the able-bodied references. The powered prosthetic joint trajectories show strong similarities to the intact joint trajectories, which also indicate that the quai-stiffness can improve gait symmetry. Fig. 8 shows the root mean squared error (RMSE) in the powered prosthetic joint trajectories relative to able-bodied trajectories and intact leg trajectories during the walking experiments. The lower RMSE of the observed kinematics with respect to the intact leg kinematics indicates that the proposed quasi-stiffness control achieves better joint kinematic symmetry. For comparison, Fig. 9 shows the observed mean kinematics and kinetics trajectories of the proposed learning quasi-stiffness control framework and the benchmark FSMIC. Fig. 10 shows the RMSE of the observed kinematics and kinetics generated by the proposed method and the benchmark FSMIC with respect to the able-bodied references. The angle and moment RMSE of the proposed method are usually smaller than the benchmark.